media%2Ffc1%2Ffc1af120-a842-45bc-b661-44
Answer #1

a.

Coefficient of Variation = Standard Deviation / return

X = 0.35/0.1 = 3.5

Y = 0.25/0.125 = 2

b.

Higher standard deviation means returns our father away from the mean and therefore more riskier. Stock X is riskier.

c.

Required rate of return = Risk Free Rate + Beta*(Market Risk Premium)

X = 6% +0.9(5%) = 10.5%

Y = 6% + 1.2(5%) = 12%

d.

For a diversified investor stock y i will be more attractive

e.

Required rate of return of portfolio with 75% X and 25% Y

Return = 75%*10.5 + 25%*12 = 10.875%

f.

Stock with the higher beta will have larger increase in required return

X = 6% +0.9(6%) = 11. 4%

Y = 6% + 1.2(6%) = 13.2%

Therefore stock Y with higher beta had a larger increase in required return